OpenStudy (lakers):

Classify the triangle based on the side lengths 4.2, 6.4 and 7.6. right acute obtuse no triangle can be formed with given side lengths

OpenStudy (zeesbrat3):

@lakers The other problem was just the same as this one. Try using the same steps we showed you to find the answer.

OpenStudy (zeesbrat3):

This is a triangle though 4.2+6.4>7.6 6.4+7.6>4.2 7.6+4.2>6.4

OpenStudy (lakers):

its acute right?

OpenStudy (zeesbrat3):

Well its not a right triangle. I would use trig to find the angle personally.

OpenStudy (lakers):

i havent took trig yet...

OpenStudy (zeesbrat3):

Oh.. What math?

OpenStudy (lakers):

this is geometry....

OpenStudy (zeesbrat3):

You havent taken trig yet?

OpenStudy (lakers):

no im goin to be a senior this fall

OpenStudy (zeesbrat3):

Oh, but do you know basic trig? Like tan,sin, and cos?

OpenStudy (lakers):

i failed geo freshmen yr so i got to retake it on this cp class

OpenStudy (lakers):

i believe i had a lil of that

OpenStudy (zeesbrat3):

tan=o/a sin=o/h cos=a/h